OpenStudy (anonymous):

Solve. 19 − 3(2t −1) =10? show work

OpenStudy (saifoo.khan):

19 - 6t + 3 = 10 -6t + 22 = 10 -6t = 10 -22 -6t = -12 Cancel the -ves. 6t = 12 t = 12/6 t = 2